A 0.31 F capacitor is desired. What area must the plates have if they are to be separated by a 3.2 mm air gap? In m2
Answer #1

Parallel plate cap
C = ????(A/d) in Farads
?? is 8.8542e-12 F/m
?? is dielectric constant (vacuum = 1)
A and d are area of plate in m
Answer #2

C = eo*A/d....

A = C*d/eo = (0.31*3.2*10^-3)/(8.85*10^-12) = 1.12*10^8 m^2